The charges and coordinates of two charged particles held fixed in an xy plane are q1 = 2.35 μC, x1 = 3.13 cm, y1 = 0.644 cm and q2 = -6.36 μC, x2 = -2.04 cm, y2 = 1.40 cm. Find the (a) magnitude and (b) direction of the electrostatic force on particle 2 due to particle 1. At what (c) x and (d) y coordinates should a third particle of charge q3 = 3.72 μC be placed such that the net electrostatic force on particle 2 due to particles 1 and 3 is zero?

(a) is 49 N
Answer #1

q1 will attract q2.

r = (3.13 + 2.04)i + (0.644 - 1.40)j = 5.17i - 0.756 j cm

|r| = sqrt(5.17^2 + 0.756 ^2) = 5.225 cm   Or 0.05225 m


F = k q1 q2 r / |r|^3


= (9 x 10^9 x 2.35 x 10^-6 x 6.36 x 10^-6) (5.17i - 0.756j) (10^-2) / (0.05225^3)


= 48.75i - 7.13 j N


magnitude = sqrt(48.75^2 + 7.13^2) = 49.27 N

(B) Direction, @ = - tan^-1(7.13 / 48.75) = 8.32 deg below +x axis.